# Combining Evidence Across Filtrations

### v3

@@ 2.4 One-to-one correspondence between p-to-e calibrators and adjusters
(denoting $1/\infty=0$). To see this, observe that for any p-to-e calibrator $\mathsf{C}:[0,1]\to[0,\infty]$, we have that $\mathsf{A}(\mathfrak{e})=\mathsf{C}(1/\mathfrak{e})$ is right-continuous, $\mathsf{A}(\infty)=\mathsf{C}(0)=\infty$, and
[…]
@@ 3.1 Combining with a p-process in a sub-filtration via $\mathfrak{p}$-lifting
We first establish that any p-process can be lifted [^2] This notion of lifting across filtrations is largely distinct from the notion of lifting in early studies of complex-valued, continuous-time martingales (Getoor and Sharpe,, 1972). See Section E for additional context. into a finer filtration without calibration. This somewhat surprising result is a direct consequence of the “equivalence lemma” (Ramdas et al.,, 2020; Howard et al.,, 2021) between time-uniform validity, random time validity, and anytime-validity of sequential tests. Specifically, given a p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ in a filtration $\mathbb{F}$, the lemma says that the following statements involving the probability of erroneously rejecting the null are equivalent, given any $\alpha\in(0,1)$ and for each $P\in\mathcal{P}$:
@@
We give a self-contained proof in Section F.1. Only the lifting direction, a $\Rightarrow$ b, is nontrivial: it states that, if any fixed upper bound on the probability of a $\mathbb{G}$-adapted event holds at all stopping times in $\mathbb{G}$, then the bound further holds at any stopping time in any finer filtration $\mathbb{F}\supseteq\mathbb{G}$. Note that it is the coarser filtration $\mathbb{G}$ that the sequence of events has to be adapted to—if $(\xi_{t})_{t\geq 0}$ were adapted to $\mathbb{F}$ but not $\mathbb{G}$, then the equivalence would no longer hold. We also remark that the analogous claim involving upper bound statements about an expectation (e.g., $\mathbb{E}_{P}[\mathfrak{e}_{\tau}]\leq 1$) is not true in general. This necessitates the use of an adjuster when lifting e-processes (as opposed to p-processes), which we will discuss in the next subsection.
@@ 3.2 Combining with an e-process in a sub-filtration via $\mathfrak{e}$-lifting
The proof of Theorem 2, included in Section F.2, consists of three steps:
@@
A useful side effect of using adjusters for $\mathfrak{e}$-lifting is that taking the running maximum of the e-process comes for free. (In Section 5, we empirically illustrate this side effect in a concrete example.) Nevertheless, as alluded to in Section 2.3, using an adjuster involves sacrificing a logarithmic amount of evidence, meaning that achieving $\mathbb{F}$-anytime-validity itself is not free. In Section 6, we formalize a sense in which this is a necessary cost; in Section G.1, we further provide an empirical study of the statistical power when using adjusters.
@@ 4.1 Simulation results under the null and alternative hypotheses
Figure 4: Unlike the conformal test martingale (solid orange), which is not $\mathbb{F}$-anytime-valid, its adjusted version via $\mathfrak{e}$-lifting (dashed green) is $\mathbb{F}$-anytime-valid. The plot shows these two processes at an $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$: the first time seeing five consecutive zeros (4). Over $10,000$ runs, the adjusted conformal test martingale has an estimated mean of $0.47\pm 0.003$ at $\tau^{\mathbb{F}}$, contrasting with the unadjusted version. The adjusted e-process can then be combined seamlessly with any other $\mathbb{F}$-anytime-valid e-process, such as the UI e-process.
@@
Figure 5: UI (dashed blue), conformal (dotted orange), and the combined (“eLift+Avg”, solid green; Corollary 2) e-processes for testing randomness in two scenarios where the binary data is sampled from non-random distributions (left: first-order Markov; right: IID with sharp changepoints). The combined e-process grows large under both alternatives, while the UI and conformal e-processes only grow large under the Markov and changepoint alternatives, respectively. Each line is averaged over 1,000 repeated runs.
Recall from Section 1.2 that we could not simply average the two e-processes $\mathfrak{e}^{\mathsf{UI}}$ and $\mathfrak{e}^{\mathsf{conf}}$ for testing randomness, as the conformal test martingale $\mathfrak{e}^{\mathsf{conf}}$ is not $\mathbb{F}$-anytime-valid. By Corollary 2, we can now combine the two e-process by lifting $\mathfrak{e}^{\mathsf{conf}}$ into $\mathbb{F}$ using an adjuster.
@@
As shown in the plots, unlike the conformal test martingale, its adjusted version is mostly no greater than 1. Over $10,000$ repeated runs, the mean of the stopped e-values for the adjusted e-process is $0.47\pm 0.003$, suggesting that the adjusted e-process meets the validity definition (1) at this $\mathbb{F}$-stopping time. This gives evidence that the adjusted e-process can be combined with the UI e-process via averaging, such that the combined version is an e-process in $\mathbb{F}$.
Next, we demonstrate how the combined e-process via $\mathfrak{e}$-lifting can leverage the statistical power of UI and conformal e-processes, which are powerful against different alternative hypotheses. In Figure 5, we compare UI, conformal, and the combined e-processes on a simulated, non-random binary data sequence. Following Corollary 2, we construct the combined e-process as $\bar{\mathfrak{e}}^{\mathsf{adj}}=(\bar{\mathfrak{e}}_{t}^{\mathsf{adj}})_{t\geq 0}$, where $\bar{\mathfrak{e}}_{0}^{\mathsf{adj}}=1$ and
@@
First, we generate data from a first-order Markov chain with $p_{0\to 1}=0.5$ and $p_{1\to 1}=0.4$, where $p_{j\to k}$ denotes the transition probability from state $j$ to state $k$, for $j,k\in\{0,1\}$. Then, in Figure 5(a), we plot the three e-processes, each averaged over 1,000 repeated samples of length $T=2,500$. The plot confirms that the UI e-process (dashed blue) grows large while the conformal test martingale (dotted orange) does not; more importantly, it shows that the combined e-process (solid green) closely tracks the fast-growing UI e-process. This is expected as the combined e-process $\bar{\mathfrak{e}}^{\mathsf{adj}}$ is effectively half of the exponentially growing UI e-process $\mathfrak{e}^{\mathsf{UI}}$.
Next, we generate data from a drifting distribution that is typically an IID $\mathsf{Ber}(0.5)$ (“usual state”), but after the first 500 steps it briefly switches to an IID $\mathsf{Ber}(0.2)$ (“unusual state”) for 100 steps, before returning to the usual state for the remainder of the sequence. The resulting data contains two changepoints over the sequence of length $T=2,500$. As mentioned earlier, conformal test martingales are known to be powerful against changepoint alternatives, unlike the UI e-process. Figure 5(b), which shows the three e-processes averaged over 1,000 repeated runs, confirms this: the conformal test martingale grows large upon entering the unusual state and then gradually decreases upon reentering the usual state. In contrast, the UI e-process remains small throughout the sequence. Because one of the two processes grows large (in the unusual state), analogous to the Markovian case, the combined e-process also grows large.
Unlike in Figure 5(a), however, in Figure 5(b) we additionally observe two peculiarities stemming from the $\mathfrak{e}$-lifting procedure. First, because the combined e-process utilizes the adjusted conformal test martingale, it grows less quickly ($\approx 10^{2}$ at $t=500$) than the conformal test martingale ($\approx 10^{4}$). In this example, the adjusted evidence still exceeds a typical threshold for a “strong” evidence (Shafer,, 2021). (When using an e-process for a level-$\alpha$ sequential test (7), setting $\alpha$ to $0.1$, $0.05$, or $0.01$ corresponds to rejecting the null when the e-value exceeds 10, 20, or 100, respectively.) Second, as a byproduct of using an adjuster, the combined e-process retains the largest evidence from the conformal test martingale over time (the running maximum). As the data returns to its usual state (past time 600), the conformal test martingale gradually decreases, while the combined e-process eventually exceeds the conformal test martingale. Thus, while adjusting the e-process involves sacrificing some amount of evidence, it also comes with the useful byproduct of being able to insure against losing evidence over time.
In Appendix G.1, we include a simulated study on the statistical power of tests induced by the combined e-process (18) against that of the simple mean without adjustment (3).
@@ 4.2 Testing the randomness of high-volatility days in financial returns
Using data from various real-world financial assets, we compute the UI e-process $\mathfrak{e}_{t}^{\mathsf{UI}}$, the adjusted conformal test martingale $\mathsf{A}_{\mathsf{mix}}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})$, and their combined average e-process $\bar{\mathfrak{e}}_{t}^{\mathsf{adj}}=\frac{1}{2}[\mathfrak{e}_{t}+\mathsf{A}_{\mathsf{mix}}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})]$, for each $t=1,\dotsc,T$ where $t$ indexes all trading days from January 1, 2019 to December 31, 2023 (5 years total). Figure 7 shows these e-processes for four financial assets: the S&P 500 index (top left), Amazon stock (top right), Nike stock (bottom left), and the US Dollar to Chinese Yuan (USD-CNY) exchange rate (bottom right).
These cases illustrate different ways in which the combined e-process, via $\mathfrak{e}$-lifting, leverages the statistical power of its components. For S&P 500 and Nike, the UI e-process grows more quickly than the adjusted conformal test martingale around March 2020 (coinciding with the global Covid-19 outbreak), suggesting that the cluster of high-volatility days was more likely Markovian than IID during this time. On the other hand, during the early-to-middle part of 2022 (coinciding with peak inflation rates in the U.S.), the UI e-process decreases, while the adjusted conformal test martingale grows large, suggesting the presence of a changepoint in volatility during this time. Across both periods, only the combined e-process enjoys the growth of either e-process; if we were running these experiments live, we would have successfully detected both types of deviations from IID. Amazon and USD-CNY present different cases, where the adjusted conformal test martingale grows either comparably to the UI e-process (Amazon in 2020) or more quickly (Amazon in 2022; USD-CNY in 2023). The combined e-process, by construction, enjoys growth in either e-process (up to a constant multiplicative factor).
While we focused on a binarized notion of volatility for simplicity, we can apply an analogous procedure to the continuous-valued residuals from a volatility model, by replacing the UI e-process with Saha and Ramdas, (2024)’s e-process (powerful against autoregressive alternatives) and using the continuous analog of Vovk, 2021b ’s conformal test martingale.
@@ 5.1 Combining e-processes for sequential independence testing
The first example comes from sequential (nonparametric) independence testing (Balsubramani and Ramdas,, 2016; Podkopaev et al.,, 2023; Podkopaev and Ramdas,, 2023; Shekhar and Ramdas,, 2024; Henzi and Law,, 2024). This represents a scenario where both e-processes are constructed in a coarsening of the data filtration. By a complete analogy to Corollary 2, we can combine these e-processes by adjusting each and then averaging them.
@@ 5.2 Evaluating and comparing $h$-step-ahead sequential forecasters
We proceed with a concrete example based on the formulations of Choe and Ramdas, (2023, appendix E) and Arnold et al., (2023, section 4). Let $h>1$ be an integer lag, and suppose that at each time step $t=1,2,\dotsc$, two forecasters $\mathsf{p}$ and $\mathsf{q}$ make respective probability forecasts $p_{t},q_{t}\in[0,1]$ for an eventual binary outcome $y_{t+h-1}\in\{0,1\}$, say, whether it will rain on the $(t+h-1)$th day. The underlying distribution for the outcome sequence is not known. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration with which both $p_{t}$ and $q_{t}$ are $\mathcal{F}_{t-1}$-measurable—that is, $\mathcal{F}_{t-1}$ represents the information available to each forecaster before making their forecast on $y_{t+h-1}$.
@@
where $I_{t}^{[k]}$ is an index set that includes every $h$th time step starting at $k$ up to $t-h+1$. The parameter $\Delta_{t}^{[k]}$ denotes the mean conditional expectation of score differences between the two forecasters up to time $t$ at offset $k$. For $h=2$, $\Delta_{t}^{[1]}$ and $\Delta_{t}^{[2]}$ represent the mean score difference on odd and even days, respectively. We condition the scores of $p_{i}$ and $q_{i}$ on $\mathcal{F}_{i-1}$ to adequately capture the information set available to the forecasters at the time of their forecasting.
@@
The $h$-step-ahead forecast comparison problem presents a unique application of $\mathfrak{e}$-lifting, as there are ways to bypass the construction of a combined e-process if one only cares about constructing a sequential test. In particular, as noted (implicitly) in prior work, we may directly utilize the “free” $\mathfrak{p}$-lifting procedure (Theorem 1) given the particular problem setup. To be clear, if our goal is to construct a combined e-process instead, then using an adjuster (25) remains the most straightforward and effective way. We elaborate on these points and include an empirical comparison of the resulting approaches, in Section G.2.

### v4

@@ 2.4 One-to-one correspondence between p-to-e calibrators and adjusters
(denoting $1/\infty=0$). To see this, observe that for any admissible p-to-e calibrator $\mathsf{C}:[0,1]\to[0,\infty]$, we have that $\mathsf{A}(\mathfrak{e})=\mathsf{C}(1/\mathfrak{e})$ is right-continuous, $\mathsf{A}(\infty)=\mathsf{C}(0)=\infty$, and
$\int_{1}^{\infty}\frac{\mathsf{A}(\mathfrak{e})}{\mathfrak{e}^{2}}d\mathfrak{e}=\int_{1}^{\infty}\frac{\mathsf{C}(1/\mathfrak{e})}{\mathfrak{e}^{2}}d\mathfrak{e}=\int_{0}^{1}\mathsf{C}(\mathfrak{p})d\mathfrak{p}=1,$ (14)
@@ 3.1 Combining with a p-process in a sub-filtration via $\mathfrak{p}$-lifting
We first establish that any p-process can be lifted [^2] This notion of lifting across filtrations is largely distinct from the notion of lifting in early studies of complex-valued, continuous-time martingales (Getoor and Sharpe,, 1972). See Section F for additional context. into a finer filtration without calibration. This somewhat surprising result is a direct consequence of the “equivalence lemma” (Ramdas et al.,, 2020; Howard et al.,, 2021) between time-uniform validity, random time validity, and anytime-validity of sequential tests. Specifically, given a p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ in a filtration $\mathbb{F}$, the lemma says that the following statements involving the probability of erroneously rejecting the null are equivalent, given any $\alpha\in(0,1)$ and for each $P\in\mathcal{P}$:
@@
We give a self-contained proof in Section G.1. Only the lifting direction, a $\Rightarrow$ b, is nontrivial: it states that, if any fixed upper bound on the probability of a $\mathbb{G}$-adapted event holds at all stopping times in $\mathbb{G}$, then the bound further holds at any stopping time in any finer filtration $\mathbb{F}\supseteq\mathbb{G}$. Note that it is the coarser filtration $\mathbb{G}$ that the sequence of events has to be adapted to—if $(\xi_{t})_{t\geq 0}$ were adapted to $\mathbb{F}$ but not $\mathbb{G}$, then the equivalence would no longer hold. We also remark that the analogous claim involving upper bound statements about an expectation (e.g., $\mathbb{E}_{P}[\mathfrak{e}_{\tau}]\leq 1$) is not true in general. This necessitates the use of an adjuster when lifting e-processes (as opposed to p-processes), which we will discuss in the next subsection.
@@ 3.2 Combining with an e-process in a sub-filtration via $\mathfrak{e}$-lifting
The proof of Theorem 2, included in Section G.2, consists of three steps:
@@
A useful side effect of using adjusters for $\mathfrak{e}$-lifting is that taking the running maximum of the e-process comes for free. (In Section 5, we empirically illustrate this side effect in a concrete example.) Nevertheless, as alluded to in Section 2.3, using an adjuster involves sacrificing a logarithmic amount of evidence, meaning that achieving $\mathbb{F}$-anytime-validity itself is not free. In Section 6, we formalize a sense in which this is a necessary cost; in Section H.1, we further provide an empirical study of the statistical power when using adjusters.
@@ 4.1 Simulation results under the null and alternative hypotheses
Figure 4: Unlike the conformal test martingale (solid orange), which is not $\mathbb{F}$-anytime-valid, its adjusted version via $\mathfrak{e}$-lifting (dashed green) is $\mathbb{F}$-anytime-valid. The plot shows these two processes at an $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$: the first time seeing five consecutive zeros (4). Over $10,000$ runs, the adjusted CTM has an estimated mean of $0.60\pm 0.001$ at $\tau^{\mathbb{F}}$, contrasting with the unadjusted version. The adjusted e-process can then be combined seamlessly with any other $\mathbb{F}$-anytime-valid e-process, such as the UI e-process.
@@
Figure 5: UI (dashed blue), conformal (dotted orange; the “simple jumper” variant), and the combined (solid green; Corollary 2) e-processes for testing randomness in two scenarios where the binary data is sampled from non-random distributions (left: first-order Markov; right: IID with sharp changepoints). The combined e-process grows large under both alternatives, while the UI and conformal e-processes only grow large under the Markov and changepoint alternatives, respectively. Each line is averaged over 1,000 repeated runs.
Recall from Section 1.2 that we could not simply average the two e-processes $\mathfrak{e}^{\mathsf{UI}}$ and $\mathfrak{e}^{\mathsf{conf}}$ for testing randomness, as the conformal test martingale (CTM) $\mathfrak{e}^{\mathsf{conf}}$ is not $\mathbb{F}$-anytime-valid. By Corollary 2, we can now combine the two e-process by lifting $\mathfrak{e}^{\mathsf{conf}}$ into $\mathbb{F}$ using an adjuster.
@@
As shown in the plots, unlike the original CTM, its adjusted version is mostly no greater than 1. Over $10,000$ repeated runs, the mean of the stopped e-values for the adjusted e-process is $0.60\pm 0.001$, suggesting that the adjusted e-process meets the validity definition (1) at this $\mathbb{F}$-stopping time. This gives evidence that the adjusted e-process can be combined with the UI e-process via averaging, such that the combined version is an e-process in $\mathbb{F}$.
Next, we show how the combined e-process via $\mathfrak{e}$-lifting can leverage the statistical power of UI and conformal e-processes, which are powerful against different alternative hypotheses. In Figure 5, we compare UI, conformal (the “simple jumper” variant; see Section B.3), and the combined e-processes on a simulated, non-random binary data sequence. Following Corollary 2, we construct the combined e-process as $\bar{\mathfrak{e}}^{\mathsf{adj}}=(\bar{\mathfrak{e}}_{t}^{\mathsf{adj}})_{t\geq 0}$, where $\bar{\mathfrak{e}}_{0}^{\mathsf{adj}}=1$ and
@@
First, we generate data from a first-order Markov chain with $p_{0\to 1}=0.5$ and $p_{1\to 1}=0.4$, where $p_{j\to k}$ denotes the transition probability from state $j$ to state $k$, for $j,k\in\{0,1\}$. Then, in Figure 5(a), we plot the three e-processes, each averaged over 1,000 repeated samples of length $T=2,500$. The plot confirms that the UI e-process (dashed blue) grows large while the CTM (dotted orange) does not; more importantly, it shows that the combined e-process (solid green) closely tracks the fast-growing UI e-process. This is expected as the combined e-process $\bar{\mathfrak{e}}^{\mathsf{adj}}$ is effectively half of the exponentially growing UI e-process $\mathfrak{e}^{\mathsf{UI}}$.
Next, we generate data from a drifting distribution that is typically an IID $\mathsf{Ber}(0.5)$ (“usual state”), but after the first 500 steps it briefly switches to an IID $\mathsf{Ber}(0.2)$ (“unusual state”) for 100 steps, before returning to the usual state for the remainder of the sequence. The resulting data contains two changepoints over the sequence of length $T=2,500$. As mentioned earlier, CTMs are known to be powerful against changepoint alternatives, unlike the UI e-process. Figure 5(b), which shows the three e-processes averaged over 1,000 repeated runs, confirms this: the CTM grows large upon entering the unusual state and then gradually decreases upon reentering the usual state. In contrast, the UI e-process remains small throughout the sequence. Because one of the two processes grows large (in the unusual state), analogous to the Markovian case, the combined e-process also grows large.
Unlike in Figure 5(a), however, in Figure 5(b) we additionally observe two peculiarities stemming from the $\mathfrak{e}$-lifting procedure. First, because the combined e-process utilizes the adjusted CTM, it grows less quickly ($\approx 10^{2}$ at $t=500$) than the CTM ($\approx 10^{4}$). In this example, the adjusted evidence still exceeds a typical threshold for a “strong” evidence (Shafer,, 2021). (When using an e-process for a level-$\alpha$ sequential test (7), setting $\alpha$ to $0.1$, $0.05$, or $0.01$ corresponds to rejecting the null when the e-value exceeds 10, 20, or 100, respectively.) Second, as a byproduct of using an adjuster, the combined e-process retains the largest evidence from the CTM over time (the running maximum). As the data returns to its usual state (past time 600), the CTM gradually decreases, while the combined e-process eventually exceeds the CTM. Thus, while adjusting the e-process involves sacrificing some amount of evidence, it also comes with the useful byproduct of being able to insure against losing evidence over time.
In Appendix H.1, we include a simulated study on the statistical power of tests induced by the combined e-process (18) against that of the simple mean without adjustment (3).
@@ 4.2 Testing the randomness of high-volatility days in financial returns
Using data from various real-world financial assets, we compute the UI e-process $\mathfrak{e}_{t}^{\mathsf{UI}}$, the adjusted simple jumper CTM $\mathsf{A}_{\mathsf{mix}}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})$, and their combined e-process $\bar{\mathfrak{e}}_{t}^{\mathsf{adj}}=\frac{1}{2}[\mathfrak{e}_{t}+\mathsf{A}_{\mathsf{mix}}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})]$, for each $t=1,\dotsc,T$ where $t$ indexes all trading days from January 1, 2019 to December 31, 2023 (5 years total). Figure 7 shows these e-processes for four financial assets: the S&P 500 index (top left), Amazon stock (top right), Nike stock (bottom left), and the US Dollar to Chinese Yuan (USD-CNY) exchange rate (bottom right).
These cases illustrate different ways in which the combined e-process, via $\mathfrak{e}$-lifting, leverages the statistical power of its components. For S&P 500 and Nike, the UI e-process grows more quickly than the adjusted CTM around March 2020 (coinciding with the global Covid-19 outbreak), suggesting that the cluster of high-volatility days was more likely Markovian than IID during this time. On the other hand, during the early-to-middle part of 2022 (coinciding with peak inflation rates in the U.S.), the UI e-process decreases, while the adjusted CTM grows large, providing evidence for the presence of a changepoint during this time. Across both periods, only the combined e-process enjoys the growth of either e-process; if we were running these experiments live, we would have successfully detected both deviations from IID. Amazon and USD-CNY present different cases, where the adjusted CTM grows either comparably to the UI e-process (Amazon in 2020) or more quickly (Amazon in 2022; USD-CNY in 2023). The combined e-process enjoys growth in either e-process (up to a constant multiplicative factor).
While we focused on a binarized notion of volatility for simplicity, we can apply an analogous procedure to the continuous-valued residuals from a volatility model, by replacing the UI e-process with Saha and Ramdas, (2024)’s e-process (powerful against autoregressive alternatives) and using the continuous analog of the CTM (Vovk, 2021b, ).
@@ 5.1 Combining e-processes for sequential independence testing
The first example is sequential (nonparametric) independence testing (Balsubramani and Ramdas,, 2016; Podkopaev et al.,, 2023; Podkopaev and Ramdas,, 2023; Shekhar and Ramdas,, 2024; Henzi and Law,, 2024). This represents a scenario where both e-processes are constructed in a coarsening of the data filtration. By a complete analogy to Corollary 2, we can combine these e-processes by adjusting each and then averaging them.
@@ 5.2 Evaluating and comparing $h$-step-ahead sequential forecasters
We proceed with a concrete example based on the formulations of Choe and Ramdas, (2023, appendix E) and Arnold et al., (2023, section 4). Let $h>1$ be an integer lag, and suppose that at each time step $t=1,2,\dotsc$, two forecasters $\mathsf{p}$ and $\mathsf{q}$ make respective probability forecasts $p_{t},q_{t}\in[0,1]$ for an eventual binary outcome $y_{t+h-1}\in\{0,1\}$, say, whether it will rain on the (t+h1)(t+h-1)th day. The underlying distribution for the outcome sequence is not known. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration with which both $p_{t}$ and $q_{t}$ are $\mathcal{F}_{t-1}$-measurable—that is, $\mathcal{F}_{t-1}$ represents the information available to each forecaster before making their forecast on $y_{t+h-1}$.
@@
where $I_{t}^{[k]}$ is an index set that includes every $h$th time step starting at $k$ up to th+1t-h+1. The parameter $\Delta_{t}^{[k]}$ denotes the mean conditional expectation of score differences between the two forecasters up to time $t$ at offset $k$. For $h=2$, $\Delta_{t}^{[1]}$ and $\Delta_{t}^{[2]}$ represent the mean score difference on odd and even days, respectively. We condition the scores of $p_{i}$ and $q_{i}$ on $\mathcal{F}_{i-1}$ to adequately capture the information set available to the forecasters at the time of their forecasting.
@@
The $h$-step-ahead forecast comparison problem presents a unique application of $\mathfrak{e}$-lifting, as there are ways to bypass the construction of a combined e-process if one only cares about constructing a sequential test. In particular, as noted (implicitly) in prior work, we may directly utilize the “free” $\mathfrak{p}$-lifting procedure (Theorem 1) given the particular problem setup. To be clear, if our goal is to construct a combined e-process instead, then using an adjuster (25) remains the most straightforward and effective way. We elaborate on these points and include an empirical comparison of the resulting approaches, in Section H.2.